Mr. and Mrs. Chaulk have three dependent children, ages 3, 6, and 9. Assume the taxable year is 2019.

  1. Compute their child credit if AGI on their joint return is $88,300.
  2. Compute their child credit if AGI on their joint return is $462,700.
  3. Compute their child credit if AGI on their joint return is $200,000 and assume that they have one non-child dependent who meets the requirements for the child credit.
Answer #1

to qualify for child credit , child should be dependent and below 17 years of age.

All three children qualifies for child credit. child credit is $2000 Per child.

Phase out limit is $400,000 for joint returns.

(1) AGI is less than 400,000 $ and all dependent children are within the age limit prescribed i.e. 17 years.

so Mr and Mrs. Chaulk are eligible for $2000 per child * 3qualifying child

= $ 6000 child Tax credit

(2) AGI is more than phase out limit of 400,000 $ so we will calculate child credit as follows:

for each 1000$ above 400,000 $50 is reduced.

=462,700-400,000

=62,700

=62700/1000 =62

=62*50$ = 3100 $

eligble credit = 6000 $ (as above ) - 3100 $

=2900 $

(3)To claim for child tax credit dependent must be son, daughter, stepson, stepdaughter, brother , sister, step brother, step sister or any descendants.

as they have one non child dependent they would not qualify for child tax credit however they can claim qualified dependent credit which is $500 per year.

so $6000 for three dependent children and 500 $ for one non child dependent

= $ 6500